Benjamin Company had the following results of operations for the past year: $ 157,600 Sales (16,000 units at $9.85) Direct maMultiple Choice Ο Increase by $4,884. Ο Decease by $4,884. Ο Increase by $26,529. Ο Increase by $4,255. Ο Increase by $3,415.
Answer #1

Answer :- Calculation of the Profit if Benjamin accept the offer :-

Direct Material and Direct labor per unit=$93,600/16,000=$5.85

Relevant Variable Overhead = $13,600 * 20% = $2,720

Relevant Variable Cost per unit = $2,720 / 16,000 = $0.17

Total Relevant Variable cost per unit = $5.85 + $0.17 = $6.02

Relevant Contribution Margin per unit = $7.17 - $6.02 = $1.15

Total Contribution = 3,700 * $1.15 = $4,255

Profit = Contribution - Fixed Cost

Profit = $4,255 - $570 - $270

Profit = $3,415

So the correct answer is option E. ( $3,415).

Working :-

Selling and administrative expenses are all fixed and will not change due to change in production. So these costs will be ignored.

If you have any query ask in comment section. If you like the answer plz rate. Thanks